# Robust Molecular Property Prediction via Densifying Scarce Labeled Data

### v2

@@ Abstract
A widely recognized limitation of molecular prediction models is their reliance on structures observed in the training data, resulting in poor generalization to out-of-distribution compounds. Yet in drug discovery, the compounds most critical for advancing research often lie beyond the training set, making the bias toward the training data particularly problematic. This mismatch introduces substantial covariate shift, under which standard deep learning models produce unstable and inaccurate predictions. Furthermore, the scarcity of labeled data—stemming from the onerous and costly nature of experimental validation—further exacerbates the difficulty of achieving reliable generalization. To address these limitations, we propose a novel meta-learning-based approach that leverages unlabeled data to interpolate between in-distribution (ID) and out-of-distribution (OOD) data, enabling the model to meta-learn how to generalize beyond the training distribution. We demonstrate significant performance gains on challenging real-world datasets with substantial covariate shift, supported by t-SNE visualizations highlighting our interpolation method.
@@ 1 Introduction
While vast collections of unlabeled molecular structures are readily available (Sterling & Irwin, 2015; Kim et al., 2021), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization (Klarner et al., 2023). Therefore, we propose a novel meta-learning based method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space. Our code can be found at https://github.com/JinA0218/drugood-densify.
@@ 2 Methodology
Scarce Data Densification with Unlabeled Data To address this, we propose a meta-learning based framework that interpolates the training distribution $\mathcal{D}_{\text{train}}$ with an exogenous distribution $\mathcal{D}_{\text{unlabeled}}$. Our objective is to leverage the cheaper and more abundant distribution $\mathcal{D}_{\text{unlabeled}}$ to densify the scarce labeled distribution $\mathcal{D}_{\text{train}}$ in a way that encourages the model to generalize robustly under covariate shift, particularly in out-of-distribution scenarios where we have no label information and therefore high uncertainty. For this, we utilize subsets of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{context}}$ and $\mathcal{D}_{\text{mvalid}}$, where $\mathcal{D}_{\text{context}}$ is a domain-informed external task distribution for interpolating to $\mathcal{D}_{\text{train}}$, and $\mathcal{D}_{\text{mvalid}}$ is a meta-validation set used to guide the interpolation function. Inspired by (Lee et al., 2024), we introduce a permutation invariant learnable set function (Zaheer et al., 2017; Lee et al., 2019) $\mu_{\lambda}$ as a mixer (interpolator), which learns to mix each point from $x_{i}\sim\mathcal{D}_{\text{train}}$ with the context points $\{c_{ij}\}_{j=1}^{m_{i}}$ in a way that densifies $\mathcal{D}_{\text{train}}$, where
@@
Model HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector $L_{1}$-Regression $1.137{\scriptstyle\pm.000}$ $0.714{\scriptstyle\pm.000}$ $1$000 $1.130{\scriptstyle\pm.000}$ $0.482{\scriptstyle\pm.000}$ $0.442{\scriptstyle\pm.000}$ $L_{2}$-Regression 0.999±.000plus-or-minus0.999.0000.$9$000 $0.723{\scriptstyle\pm.000}$ $1.495{\scriptstyle\pm.000}$ $1.143{\scriptstyle\pm.000}$ $0.498{\scriptstyle\pm.000}$ $0.436{\scriptstyle\pm.000}$ Random Forest $0.815{\scriptstyle\pm.009}$ $0.834{\scriptstyle\pm.010}$ $1.473{\scriptstyle\pm.008}$ $1.461{\scriptstyle\pm.012}$ $0.458{\scriptstyle\pm.002}$ $0.438{\scriptstyle\pm.002}$ MLP $0.768{\scriptstyle\pm.014}$ $2.118{\scriptstyle\pm.015}$ $1.393{\scriptstyle\pm.024}$ $1.094{\scriptstyle\pm.029}$ $0.443{\scriptstyle\pm.007}$ $0.399{\scriptstyle\pm.006}$ Mixup $0.764{\scriptstyle\pm 0.008}$ $0.691{\scriptstyle\pm 0.022}$ $1.439{\scriptstyle\pm 0.021}$ $1.212{\scriptstyle\pm 0.012}$ $0.481{\scriptstyle\pm 0.002}$ $0.479{\scriptstyle\pm 0.003}$ Mixup (w/ outlier exposure) $0.748{\scriptstyle\pm 0.01}$ $0.677{\scriptstyle\pm 0.015}$ $1.384{\scriptstyle\pm 0.012}$ $1.224{\scriptstyle\pm 0.016}$ $0.442{\scriptstyle\pm 0.005}$ $0.443{\scriptstyle\pm 0.005}$ Manifold Mixup $0.88{\scriptstyle\pm 0.023}$ $0.898{\scriptstyle\pm 0.022}$ $1.414{\scriptstyle\pm 0.021}$ $1.367{\scriptstyle\pm 0.04}$ $0.432{\scriptstyle\pm 0.005}$ $0.499{\scriptstyle\pm 0.013}$ Manifold Mixup (w/ bilevel) $0.484{\scriptstyle\pm 0.011}$ $0.804{\scriptstyle\pm 0.086}$ $1.19{\scriptstyle\pm 0.05}$ $1.217{\scriptstyle\pm 0.068}$ $0.43{\scriptstyle\pm 0.013}$ $0.536{\scriptstyle\pm 0.039}$ Q-SAVI $0.682{\scriptstyle\pm.019}$ $0.664{\scriptstyle\pm.028}$ $1.332{\scriptstyle\pm.017}$ $1.028{\scriptstyle\pm.027}$ $0.436{\scriptstyle\pm.007}$ $0.387{\scriptstyle\pm.012}$ Ours (Deepsets) $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $\mathbf{0.984\pm 0.018}$ $\mathbf{0.963\pm 0.017}$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
@@ 3 Experiments
[…]
@@ 5 Conclusion
To overcome the challenge of test-time covariate shift, we propose a novel meta-learning based method that densifies the training distribution using domain-informed unlabeled datasets via a learnable set function. During training, we utilize bilevel optimization which leverages noisy unlabeled data to guide our model towards robustness under covariate shift. We validate our method on challenging real-world molecular property prediction with large covariate shifts, and visually demonstrate the effectiveness of interpolation method which shows well defined separation between data distributions. Our code is available at https://github.com/JinA0218/drugood-densify.

### v3

@@ Abstract
A widely recognized limitation of molecular prediction models is their reliance on structures observed in the training data, resulting in poor generalization to out-of-distribution compounds. Yet in drug discovery, the compounds most critical for advancing research often lie beyond the training set, making the bias toward the training data particularly problematic. This mismatch introduces substantial covariate shift, under which standard deep learning models produce unstable and inaccurate predictions. Furthermore, the scarcity of labeled data—stemming from the onerous and costly nature of experimental validation—further exacerbates the difficulty of achieving reliable generalization. To address these limitations, we propose a novel bilevel optimization approach that leverages unlabeled data to interpolate between in-distribution (ID) and out-of-distribution (OOD) data, enabling the model to learn how to generalize beyond the training distribution. We demonstrate significant performance gains on challenging real-world datasets with substantial covariate shift, supported by t-SNE visualizations highlighting our interpolation method.
@@ 1 Introduction
While vast collections of unlabeled molecular structures are readily available (Sterling & Irwin, 2015; Kim et al., 2021), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization (Klarner et al., 2023). Therefore, we propose a novel bilevel optimization method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space. Our code can be found at https://github.com/JinA0218/drugood-densify.
@@ 2 Methodology
Scarce Data Densification with Unlabeled Data To address this, we propose a bilevel optimization framework that interpolates the training distribution $\mathcal{D}_{\text{train}}$ with an exogenous distribution $\mathcal{D}_{\text{unlabeled}}$. Our objective is to leverage the cheaper and more abundant distribution $\mathcal{D}_{\text{unlabeled}}$ to densify the scarce labeled distribution $\mathcal{D}_{\text{train}}$ in a way that encourages the model to generalize robustly under covariate shift, particularly in out-of-distribution scenarios where we have no label information and therefore high uncertainty. For this, we utilize subsets of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{context}}$ and $\mathcal{D}_{\text{mvalid}}$, where $\mathcal{D}_{\text{context}}$ is a domain-informed external task distribution for interpolating to $\mathcal{D}_{\text{train}}$, and $\mathcal{D}_{\text{mvalid}}$ is a meta-validation set used to guide the interpolation function. Inspired by (Lee et al., 2024), we introduce a permutation invariant learnable set function (Zaheer et al., 2017; Lee et al., 2019) $\mu_{\lambda}$ as a mixer (interpolator), which learns to mix each point from $x_{i}\sim\mathcal{D}_{\text{train}}$ with the context points $\{c_{ij}\}_{j=1}^{m_{i}}$ in a way that densifies $\mathcal{D}_{\text{train}}$, where
@@
Model HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector $L_{1}$-Regression $1.137{\scriptstyle\pm.000}$ $0.714{\scriptstyle\pm.000}$ $611{\scriptstyle\pm.000} 1.1$$130{\scriptstyle\pm.000}$ $0.482{\scriptstyle\pm.000}$ $0.442{\scriptstyle\pm.000}$ $L_{2}$-Regression $0.999{\scriptstyle\pm.000}$ $0.723{\scriptstyle\pm.000}$ $495{\scriptstyle\pm.000} 1.1$$143{\scriptstyle\pm.000}$ $0.498{\scriptstyle\pm.000}$ $0.436{\scriptstyle\pm.000}$ Random Forest $0.815{\scriptstyle\pm.009}$ $0.834{\scriptstyle\pm.010}$ $1.473{\scriptstyle\pm.008}$ $1.461{\scriptstyle\pm.012}$ $0.458{\scriptstyle\pm.002}$ $0.438{\scriptstyle\pm.002}$ MLP $0.768{\scriptstyle\pm.014}$ $2.118{\scriptstyle\pm.015}$ $1.393{\scriptstyle\pm.024}$ $1.094{\scriptstyle\pm.029}$ $0.443{\scriptstyle\pm.007}$ $0.399{\scriptstyle\pm.006}$ Mixup $0.764{\scriptstyle\pm 0.008}$ $0.691{\scriptstyle\pm 0.022}$ $1.439{\scriptstyle\pm 0.021}$ $1.212{\scriptstyle\pm 0.012}$ $0.481{\scriptstyle\pm 0.002}$ $0.479{\scriptstyle\pm 0.003}$ Mixup (w/ outlier exposure) $0.748{\scriptstyle\pm 0.01}$ $0.677{\scriptstyle\pm 0.015}$ $1.384{\scriptstyle\pm 0.012}$ $1.224{\scriptstyle\pm 0.016}$ $0.442{\scriptstyle\pm 0.005}$ $0.443{\scriptstyle\pm 0.005}$ Manifold Mixup $0.88{\scriptstyle\pm 0.023}$ $0.898{\scriptstyle\pm 0.022}$ $1.414{\scriptstyle\pm 0.021}$ $1.367{\scriptstyle\pm 0.04}$ $0.432{\scriptstyle\pm 0.005}$ $0.499{\scriptstyle\pm 0.013}$ Manifold Mixup (w/ bilevel) $0.484{\scriptstyle\pm 0.011}$ $0.804{\scriptstyle\pm 0.086}$ $1.19{\scriptstyle\pm 0.05}$ $1.217{\scriptstyle\pm 0.068}$ $0.43{\scriptstyle\pm 0.013}$ $0.536{\scriptstyle\pm 0.039}$ Q-SAVI $0.682{\scriptstyle\pm.019}$ $0.664{\scriptstyle\pm.028}$ $1.332{\scriptstyle\pm.017}$ $1.028{\scriptstyle\pm.027}$ $0.436{\scriptstyle\pm.007}$ $0.387{\scriptstyle\pm.012}$ Ours (Deepsets) $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $\mathbf{0.984\pm 0.018}$ $\mathbf{0.963\pm 0.017}$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
@@ 3 Experiments
Furthermore, we present a series of ablation studies in Table 2 to analyze the contributions of each component of our method (Implementation details can be found in Section A.2). Specifically, to validate the effectiveness of $\mathcal{D}_{\text{context}}$ and our bilevel optimization (Lorraine et al., 2019) under $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$, we compare our approach against the following variants: (1) using $\mathcal{D}_{\text{context}}$ without bilevel optimization, (2) excluding both $\mathcal{D}_{\text{context}}$ and bilevel optimization, and (3) using the real label for $y_{i,k}^{(\text{mvalid})}$. The results indicate that, in order to achieve robust generalization across all datasets, it is crucial to leverage bilevel optimization with $\mathcal{D}_{\text{context}}$. Especially, the dataset HIVPROT, which has the highest degree of label and covariate shift compared to others (Klarner et al., 2023) (Table 3), showed a detrimental performance drop ($62.6\%\text{–}67.3\%$ for the count vector, and $63.8\%\text{–}65.3\%$ for the bit vector) in settings (1,2). Moreover, using $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$ in $L_{V}$ for training the set function $\mu_{\lambda}$ resulted in comparable or better performance than using the true label associated with $x_{i,k}^{(\text{mvalid})}$, validating our choice of $L_{V}$ and its contribution to improved robustness under covariate shift.
@@ 5 Conclusion
To overcome the challenge of test-time covariate shift, we propose a novel bilevel optimization method that densifies the training distribution using domain-informed unlabeled datasets via a learnable set function. During training, we utilize bilevel optimization which leverages noisy unlabeled data to guide our model towards robustness under covariate shift. We validate our method on challenging real-world molecular property prediction with large covariate shifts, and visually demonstrate the effectiveness of interpolation method which shows well defined separation between data distributions. Our code is available at https://github.com/JinA0218/drugood-densify.